If the demand equation is P = 30-2Q and the marginal cost of the provided service is $2, how much is a maximum up front fee for each member?
Answer #1

Maximum front fee can be determined at a point where marginal revenue equals marginal cost.

Total revenue = price * quantity

= (30-2Q)Q

= 30Q-2Q^2

Marginal revenue = 30-4Q

MR= MC

30-4Q= 2

28= 4Q

Q = 28/4

Q = 7

P = 30-2Q

= 30-2(7)

= 16

Thus, maximum up front fee charged from each member is $16.